OpenStudy (anonymous):

Solve equation 5(b+2)/2 = 5b-35/4

OpenStudy (anonymous):

\[\frac{ 5(b+2) }{ 2 }=\frac{ 5b-35 }{ 4 }\]\[\frac{ 5b+10 }{ 2 }=\frac{ 5b-35 }{ 4 }\]\[4(5b+10)=2(5b-35)\]\[20b+40=10b-75\]\[20b+40-40-10b=10b-75-10b-40\]\[10b=-115\]\[b=\frac{ -115 }{ 10 }\]

OpenStudy (anonymous):

you can simplify the fraction or change it into decimal

OpenStudy (anonymous):

I tried it n my computer is not taking those answers

OpenStudy (anonymous):

i did it manually

OpenStudy (anonymous):

This problem is on my webassign homework, n I submitted the answer n it said it was wrong.

OpenStudy (anonymous):

I tried it n my computer is not taking those answers

OpenStudy (anonymous):

is the question true?

OpenStudy (anonymous):

i mean mine

OpenStudy (anonymous):

I put that also, it showed wrong